OpenStudy (anonymous):

Simplify: -2(-3+2) - 3(3-4)^3 - 9^2 please show work

OpenStudy (anonymous):

First solve the numbers inside the quotes... so for (-3+2), you get -1. Then solve the powers (assuming you know simple multiplication) Then you do the multiplications (i.e.) -2(whatever is inside) Then you subtract or add. Its all about orders of operations

OpenStudy (anonymous):

You cannot distribute the -3 into the 3 and -4. You have to solve the power first, which means you have to solve the inside of the parenthesis first.

OpenStudy (anonymous):

-2(-3+2) - 3(3-4)^3 - 9^2 = -2(-1) - 3(-1)^3 - 81 = 2 + 3 - 81 = -76

OpenStudy (anonymous):

thank you, your explaination helped a lot